OpenStudy (anonymous):

what are the equivalent weights of glucose and co2 for equation 6CO2 + 12H2O → C6H12O6 + 6O2 +6H2O

OpenStudy (chmvijay):

equivalent weight = molar mass of glucose /no electrons involved in the following reaction

OpenStudy (anonymous):

no it's not,, the answer is 7.5 and 11 respectively but how i could not understand

OpenStudy (cwrw238):

the relative molecukar mass of glucose is 72 + 12 + 86 = 180 and of the 6 molecules of CO2 = 6*44 = 264 180 / 264 = 7.5 /11